Three iron pieces of mass m,m and 2m placed on the co-ordinates of triangle (2.5,1.5) ,(3.5,1.5) and (3,3) respectively. Find the center of mass?

Grade:11

2 Answers

Neethushree
7 years ago
We know that Com=m`x` + m"x"/m`+m"So2.5*m +3.5*m +3*2m/m+ m+ 2m =12/4=3= x-coordinate $ 1.5*m+ 1.5*m+ 3*2m/ m+m+2m=9/4=y-coordinate.Hence, com= (3,9/4).
snehasish
7 years ago
Let coordinates of cm=(X,Y). X=((∑mx)/(∑m))  similarly Y.  Now X=((2.5m+3.5m+3*2m)/(m+m+2m)) =3   and  Y =((1.5m+1.5m+3*2m)/(m+m+2m))=(9/4)